# RAG-Enabled Intent Reasoning for Application-Network Interaction

### v1

@@ Abstract
Index Terms:
Intent-based network, network automation, large language models, retrieval augmented generation.I Introduction
The plethora of evolving digital applications and services such as immersive multimedia and connected vehicles drives network operators toward automating network operations and management. IBN has recently been introduced as a promising solution to automate networks [1]. IBN relies on receiving application/service demand in an abstract and high-level description language without specifying how to achieve the desired outcome, which is called intent. IBN is responsible for translating the received intent into low-level configuration language to be executed on network functions and devices. This gives high freedom to network operators and service providers to manage and optimize their network resources. However, translating the high-level expressed intents to low-level network configuration language is a challenging problem since various applications interact with the network, each with its specialized needs and domain language. Traditional translation methods rely on creating an ontology-based language for each application to facilitate intent translation, which lacks technical expertise and is neither practical nor scalable. To tackle this problem, various efforts from both industry and academia have been made to introduce artificial intelligence (AI) and machine reasoning (MR) to mobile networks. AI and MR have powerful capabilities to interpret different intent languages and deploy translated intents into network configurations through compilers.
@@
To solve the application-network interaction problem, in this paper, we propose a framework that utilizes application domain context, RAG, LLM model with reasoning capabilities, and few-shot learning to reason over high-level application language intents (i.e., technical and non-technical) and translate them to low-level network configuration language. RAG technique is considered a class of LLM applications that use external data to augment the LLM’s context. It helps augment knowledge-intensive tasks and ensures LLMs have up-to-date knowledge. LLM models with reasoning capabilities can provide accurate translation and reasoning over the domain language of different applications. Few-shot learning is a framework that trains the LLM through a very small number of examples via a prompt. It overcomes the hallucination issue and improves the LLMs’ responses without retraining or fine-tuning, as retraining or fine-tuning the LLM is costly and consumes a lot of resources. Our main contribution can be summarized as follows:
@@
We propose a context-aware AI framework that acts as a functionality in the vertical application layer able to interpret different languages from various applications and translate them to a common language understood by the network service provider. The framework relies on the application domain context, RAG model, reasoning LLM model, and few-shot learning to preserve data privacy, provide an accurate translation, and avoid hallucination in the translated intents.
@@ II-A Problem Statement
II-B Proposed Framework
@@ III-B Evaluation
The experiment is conducted on an intent interpretation use case, where a set of network user applications express their intents in their own domain’s language to the network. Each user application generates an expectation/demand for the network in a human language. The network has to interpret and translate the application intents to structured network intents through the proposed framework. To evaluate the performance of intent-RAG and vanilla-RAG frameworks for our use case, we use both human evaluation and the RAGAS evaluator [15]. The dataset used for evaluation includes both technical and non-technical user intents and is available in the following link 222https://github.com/SalwaMostafa/RAG-Enabled-Intent-Reasoning-for-Application-Network-Interaction/blob/main/dataset. The human evaluation is presented in Table I, where the ground truth is extracted manually from the technical documents.

### v2

@@ Abstract
I Introduction
The plethora of evolving digital applications and services such as immersive multimedia and connected vehicles drives network operators toward automating network operations and management. Intent-based network (IBN) has recently been introduced as a promising solution to automate networks [11]. IBN relies on receiving application/service demand in an abstract and high-level description language without specifying how to achieve the desired outcome, which is called intent. IBN is responsible for translating the received intent into low-level configuration language to be executed on network functions and devices. This gives high freedom to network operators and service providers to manage and optimize their network resources. However, translating the high-level expressed intents to low-level network configuration language is a challenging problem since various applications interact with the network, each with its specialized needs and domain language. Traditional translation methods rely on creating an ontology-based language for each application to facilitate intent translation, which lacks technical expertise and is neither practical nor scalable. To tackle this problem, various efforts from both industry and academia have been made to introduce artificial intelligence (AI) and machine reasoning (MR) to mobile networks. AI and MR have powerful capabilities to interpret different intent languages and deploy translated intents into network configurations through compilers.
@@ I Introduction
To solve the application-network interaction problem, in this paper, we propose a framework that utilizes application domain context, RAG, LLM model with reasoning capabilities, and few-shot learning to reason over high-level application language intents (i.e., technical and non-technical) and translate them to low-level network configuration language. RAG technique is considered a class of LLM applications that use external data to augment the LLM’s context. It helps augment knowledge-intensive tasks and ensures LLMs have up-to-date knowledge. LLM models with reasoning capabilities can provide accurate translation and reasoning over the domain language of different applications. Few-shot learning [3] is a framework that trains the LLM through a very small number of examples via a prompt. It overcomes the hallucination issue and improves the LLMs’ responses without retraining or fine-tuning, as retraining or fine-tuning the LLM is costly and consumes a lot of resources. Our main contribution can be summarized as follows:
@@
We propose a context-aware AI framework that acts as a functionality in the vertical application layer able to interpret different languages from various applications and translate them to a common language understood by the network service provider. The framework relies on the application domain context, RAG model, reasoning LLM model, and few-shot learning to preserve data privacy, provide an accurate translation, and reduce hallucination in the translated intents.
@@ II-A Problem Statement
II-B Problem Formulation
We consider the problem of translating high-level, application-specific intents into structured network intents that can be directly interpreted by the network. Let denote the intent space containing natural language intents expressed by users or applications, and let denote the structured network intent space defined according to standardized information models (e.g., 3GPP, TMF). The network knowledge is stored in an external knowledge base , which contains domain-specific technical information extracted from standards, configuration manuals, and service catalogs. The translation task can be formulated as:
where denotes the conditional probability of generating a structured network intent given the natural language intent and the knowledge base .
Given an input intent , the goal of the proposed framework is to infer a structured network intent conditioned on both the intent and the relevant knowledge .
The proposed framework consists of two major stages:
Intent Refinement: A mapping , where is a well-defined intent space obtained through ambiguity resolution and context enrichment.
Structured Intent Creator: A mapping , parameterized by (e.g., the language model and prompting strategy), which generates the structured intent.
The optimal structured intent is obtained by minimizing a translation loss function between the generated intent and the ground-truth structured intent :
where is the training or evaluation dataset, and may include semantic similarity, schema conformity, or faithfulness scores.
II-C Proposed Framework
@@ III-B Evaluation
The experiment is conducted on an intent interpretation use case, where a set of network user applications express their intents in their own domain’s language to the network. Each user application generates an expectation/demand for the network in a human language. The network has to interpret and translate the application intents to structured network intents through the proposed framework. To evaluate the performance of intent-RAG and vanilla-RAG frameworks for our use case, we use both human evaluation and the RAGAS evaluator [6]. The dataset used for evaluation includes both technical and non-technical user intents and is available in [15]. The human evaluation is presented in Table I, where the ground truth is extracted manually from the technical documents.